# GTPC-SSCD: Gate-guided Two-level Perturbation Consistency-based Semi-Supervised Change Detection thanks: This work was supported in part by the Scientific Research Program of Tianjin Municipal Education Commission under Grant 2023KJ232. (Corresponding author: Rui Huang.)

### v1

@@ Abstract
Semi-supervised change detection (SSCD) employs partially labeled data and a substantial amount of unlabeled data to identify differences between images captured in the same geographic area but at different times. However, existing consistency regularization-based SSCD methods only implement perturbations at a single level and can not exploit the full potential of unlabeled data. In this paper, we introduce a novel Gate-guided Two-level Perturbation Consistency regularization-based SSCD method (GTPC-SSCD), which simultaneously maintains strong-to-weak consistency at the image level and perturbation consistency at the feature level, thus effectively utilizing the unlabeled data. Moreover, a gate module is designed to evaluate the training complexity of different samples and determine the necessity of performing feature perturbations on each sample. This differential treatment enables the network to more effectively explore the potential of unlabeled data. Extensive experiments conducted on six public remote sensing change detection datasets demonstrate the superiority of our method over seven state-of-the-art SSCD methods.
@@
Change detection (CD), consistency regularization, remote sensing, gate machine, semi-supervised learning
@@ I Introduction
Semi-supervised change detection (SSCD) aims to identify pixel-level changes between two images taken from the same scene at different times, using a limited amount of labeled data and a large amount of unlabeled data  [1, 2, 3]. It has wide applications in various fields, including natural resource monitoring and utilization  [4, 5], disaster monitoring and assessment [6], urban management and development [7, 8].
Semi-supervised methods can be divided into three categories: adversarial learning-based methods, pseudo-label-based methods, and consistency regularization-based methods. AdvEnt [9], SemiCDNet [10], and SALCD [11] are typical adversarial learning-based methods that employ alternative optimization strategies to improve the representation learning of their respective models. Pseudo-label-based methods, such as RC-CD [12], SemiSiROC [13], and DCF-CFe [14], emphasize the improvement of pseudo-label quality and apply some techniques to enhance feature distinctiveness. Consistency regularization-based methods, including SemiCD [15], SemiSANet [16], Semi-LCD [17], SemiBuildingChange [18], and SemiPTCD [19], enforce that perturbed images or features produce identical outputs as the original inputs [20].
Recently, the consistency regularization-based framework have been increasingly drawn to SSCD methods for its simplicity and stability. However, current methods often modify network architectures or introduce additional information based on image-level weak-to-strong consistency. For example, SemiSANet [16] constructs a siamese nested UNet with graph attention, while MTCNet [21] adds extra segmentation labels to change labels. ST-RCL [22] combines self-training and consistency learning for SSCD. Concurrently, feature-level consistency regularization remains relatively simplistic and underdeveloped, such as UniMatch’s implementation of a dropout to enforce feature consistency [23]. Thus, there exists significant potential in integrating image-level and feature-level consistency regularization to further enhance the performance and robustness of SSCD.
@@
Figure 1: Motivation: Comparison of SSCD with different perturbations by $5\%$ labeled training data. Sup-only denotes that our method only be trained by labeled training data. Feature: Feature-level perturbation consistency method. Image: Image-level strong-to-weak consistency method. Different colors are used for a better view, i.e., white for true positive, black for true negative, red for false positive, and green for false negative.
@@
Our research shows that employing a simple SSCD architecture and applying perturbations either at the image or feature level does not yield satisfactory performance. Fig. 1 presents the results of SSCD networks employing perturbations at different perspectives with a ratio of $5\%$ labeled data and $95\%$ unlabeled data on two CD datasets [24, 10]. Although perturbations contribute to enhancing model robustness, relying solely on image perturbations may affect pixel integrity and segmentation, and exclusive dependence on feature perturbations might struggle with image diversity.
@@
In this paper, we explore consistency regularization at a broader scope and propose a Gate-guided Two-level Perturbation Consistency regularization-based SSCD method (GTPC-SSCD). By learning both image-level strong-to-weak consistency and feature-level perturbation consistency, our method can benefit from a more thorough comprehension of the data, leading to improved robustness and discriminability. Furthermore, performing the same treatment on all samples constrains the potential effectiveness of perturbation, particularly evident in remote sensing change detection scenarios. To tackle this issue, we develop a gate module to distinguish the difficulty of training samples and determine the application of feature perturbation accordingly. As presented in Fig. 1, the performance of our method is superior to other three methods. Our contributions are summarized as follows:
@@
  • We separately analyze image and feature perturbation implementations in SSCD, and meticulously combine them with optimized strategies to improve performance.
@@
  • We introduce a gate-guided two-level perturbation consistency regularization-based SSCD method, utilizing a gate mechanism to assess the difficulty of samples and decide whether to apply feature perturbations.
@@
  • We conduct extensive experiments on six public CD datasets. Our method exhibits higher accuracy compared to other seven state-of-the-art SSCD methods.
@@ II Methodology
Semi-supervised change detection (SSCD) utilizes a limited amount of labeled data and a large amount of unlabeled data to train a change detection network for accurate change map generation. The training set consists of two subsets, a labeled set and an unlabeled set. The labeled set can be represented as $\mathcal{D}_{l}=\{{(\mathbf{X}^{l}_{A,i},\mathbf{X}^{l}_{\mathrm{B},i}),\mathbf{Y}^{l}_{i}}\}^{M}_{i=1}$, where $(\mathbf{X}^{l}_{A,i},\mathbf{X}^{l}_{B,i})$ denotes the $i$-th labeled image pair, $\mathbf{X}^{l}_{A,i}$ is the pre-change image, $\mathbf{X}^{l}_{B,i}$ is the post-change image, and $\mathbf{Y}^{l}_{i}$ is the corresponding ground truth. Let $\mathcal{D}_{u}=\{{(\mathbf{X}^{u}_{A,i},\mathbf{X}^{u}_{B,i})}\}^{N}_{i=1}$ denotes the unlabeled set. $(\mathbf{X}^{u}_{A,i},\mathbf{X}^{u}_{B,i})$ is the $i$-th unlabeled image pair. $M$ and $N$ indicate the number of labeled and unlabeled image pairs, respectively. In most cases, we have $N>>M$. In the following sections, we will introduce the implementation details of the proposed change detection network.
@@
II-A Our consistency regularization-based SSCD method
Our SSCD method consists of a supervised training part and an unsupervised training part.
In the supervised training part, we use the labeled set $\mathcal{D}_{l}$ to train the CD network $\phi$. The network processes weakly augmented image pairs to generate change map $\mathbf{P}^{l}$. We apply standard cross-entropy (CE) loss as supervision. The loss of the supervised training part is:
$\mathcal{L}_{s}=\mathcal{L}_{CE}(\mathbf{P}^{l},\mathbf{Y}^{l}).$ (1)
The unsupervised part is illustrated in Fig. 2, we apply two-level perturbations to train $\phi$ on the unlabeled set $\mathcal{D}_{u}$. Specifically, we exploit strong-to-weak consistency at the image level and perturbation consistency at the feature level. The former enhances the model’s robustness to various data transformations, while the latter improves the model’s generalization ability and resistance to interference.
Strong-to-weak consistency learning. Drawing inspiration from ConMatch [25] and ReMixMatch [26], we increase the number of strong augmentation branches and ensure the strong-to-weak consistency, fully maximizing the benefits of strong augmentation. Specifically, we apply two independent strong augmentations on $(\mathbf{X}^{uw}_{A},\mathbf{X}^{uw}_{B})$, resulting in $(\mathbf{X}^{us_{1}}_{A},\mathbf{X}^{us_{1}}_{B})$ and $(\mathbf{X}^{us_{2}}_{A},\mathbf{X}^{us_{2}}_{B})$. They are fed into $\phi$ and generate three change maps $\mathbf{P}^{uw}$, $\mathbf{P}^{us_{1}}$ and $\mathbf{P}^{us_{2}}$.
@@ II-A Our consistency regularization-based SSCD method
$\hat{\mathbf{Y}}^{uw}=\begin{cases}1,&if\ \mathbf{P}^{uw}>\tau\\ 0,&else\end{cases}\vspace{-2pt}$ (2)
where $\tau$ = 0.95 is a confidence threshold. The image-level consistency loss of the unsupervised part is as:
$\mathcal{L}_{ui}=\frac{1}{2}(\mathcal{L}_{CE}(\mathbf{P}^{us_{1}},\hat{\mathbf{Y}}^{uw})+\mathcal{L}_{CE}(\mathbf{P}^{us_{2}},\hat{\mathbf{Y}}^{uw})).\vspace{-2pt}$ (3)
In our experiments, weak augmentations consist of random resizing, cropping, and random horizontal flipping. Strong augmentations include random color jittering, Gaussian blur, and CutMix [27].
Perturbation consistency learning. We introduce feature-level consistency by applying several different perturbations on change features and enforcing the perturbation consistency. The process of generating change maps $\mathbf{P}^{uf}_{1:K}$ is depicted as:
$\mathbf{P}^{uf}_{1:K}=f^{{}^{\prime}}_{d,1:K}(\mathbf{D}^{u}_{4},\mathbf{D}^{u^{\prime}}_{1,1:K}),\vspace{-2pt}$ (4)
where $\mathbf{D}^{u^{\prime}}_{1,1:K}=\{\mathbf{D}^{u^{\prime}}_{1,k}\}^{K}_{k=1}$, represents the difference features after perturbation. $f^{{}^{\prime}}_{d,1:K}=\{f^{{}^{\prime}}_{d,k}\}^{K}_{k=1}$, denotes the auxiliary decoders. $\mathbf{P}^{uf}_{1:K}=\{\mathbf{P}^{uf}_{k}\}^{K}_{k=1}$. The feature-level consistency loss $\mathcal{L}_{uf}$ is expressed as:
$\mathcal{L}_{uf}=\frac{1}{K}\sum^{K}_{k=1}\mathcal{L}_{CE}(\mathbf{P}^{uf}_{k},\hat{\mathbf{Y}}^{uw}),\vspace{-2pt}$ (5)
where $K$ is the total number of the auxiliary decoders. We apply seven types of feature perturbations [28], i.e., feature noise, feature dropout, object masking, context masking, guided cutout, Intermediate VAT [29], and random dropout.
Hardness analysis based gate mechanism. We add an additional decoder to produce $\mathbf{P}^{uw^{\prime}}_{i}$ and conduct an Intersection over Union (IoU) comparison with $\mathbf{P}^{uw}_{i}$ to assess the complexity of each sample. Utilizing a gate mechanism, we select challenging samples and apply various perturbations on $\mathbf{D}^{u}_{1}$ of these selected samples to obtain perturbed features. The process can be formulated as:
$iou_{i}=\mathrm{IoU}(\mathbf{P}^{uw^{\prime}}_{i},\mathbf{P}^{uw}_{i}),\vspace{-1pt}$ (6)
@@
where $iou_{i}$ represents the IoU score of the $i$-th sample, $\mathrm{Mid}(iou)$ is the median IoU score. $\mathrm{FP}_{1:K}(\cdot)$ denotes $K$ different perturbations to the input features, respectively.
Loss function. The total loss consists of the supervised loss $\mathcal{L}_{s}$, the image-level consistency loss $\mathcal{L}_{ui}$ and the feature-level consistency loss $\mathcal{L}_{uf}$. It can be expressed as:
@@ II-B1 Change Detection Network
Our CD network $\phi$ consists of two components: a difference feature generator $f_{g}$ and a decoder $f_{d}$. The process of generating change map $\mathbf{P}$ is simply expressed by:
@@ II-B2 Super-parameters
We use PyTorch to conduct experiments on an NVIDIA RTX2080Ti GPU. Our model adopts the SGD optimizer with learning rate of 0.02, momentum of 0.9, and weight decay of 1e-4, respectively. All models are trained for 80 epochs with batch size of 4 for both labeled and unlabeled data.
III Experiments
III-A Experimental Setups
Baselines. We compare the proposed method with seven existing state-of-the-art methods, including AdvEnt [9], s4GAN [32], SemiCDNet [10], SemiCD [15], RC-CD [12], SemiPTCD [19], and UniMatch [23]. Sup-only refers to our method that exclusively utilizes labeled data for training.
Datasets. We conduct experiments on six widely used remote sensing CD datasets, namely WHU-CD [24], LEVIR-CD [7], BCD [24], GZ-CD [10], EGY-BCD [33], and CDD [34]. Table I details our datasets, including number of image pairs, image size, train/val/test splits, spatial resolution, and change types. The first five CD datasets mainly focus on building changes, while the last CD dataset consists of multiple types of changes, including buildings, vehicles, road expansions, and more. All images are cropped into non-overlapping patches of size $256\times 256$, which are then divided into training, validation, and test sets. The training set is further divided into labeled and unlabeled data with the following ratios: $[5\%,95\%]$, $[10\%,90\%]$, $[20\%,80\%]$, $[40\%,60\%]$.

### v2

@@ Abstract
Semi-supervised change detection (SSCD) utilizes partially labeled data and abundant unlabeled data to detect differences between multi-temporal remote sensing images. The mainstream SSCD methods based on consistency regularization have limitations. They perform perturbations mainly at a single level, restricting the utilization of unlabeled data and failing to fully tap its potential. In this paper, we introduce a novel Gate-guided Two-level Perturbation Consistency regularization-based SSCD method (GTPC-SSCD). It simultaneously maintains strong-to-weak consistency at the image level and perturbation consistency at the feature level, enhancing the utilization efficiency of unlabeled data. Moreover, we develop a hardness analysis-based gating mechanism to assess the training complexity of different samples and determine the necessity of performing feature perturbations for each sample. Through this differential treatment, the network can explore the potential of unlabeled data more efficiently. Extensive experiments conducted on six benchmark CD datasets demonstrate the superiority of our GTPC-SSCD over seven state-of-the-art methods.
@@
Change detection (CD), consistency regularization, remote sensing, gating mechanism, semi-supervised learning
@@ I Introduction
Semi-supervised change detection (SSCD) aims to identify pixel-level changes between two images taken from the same scene at different times, using a limited amount of labeled data and a large amount of unlabeled data [1]. It has wide applications in various fields, including natural resource monitoring and utilization, disaster monitoring and assessment, urban management and development [2, 3, 4].
@@
Figure 1: Motivation analysis of SSCD with different perturbation variants under $5\%$ labeled training data on WHU-CD and GZ-CD datasets. Upper part: Visual cases of different variants. Here, white for TP, black for TN, red for FP, and green for FN. Lower part: Performance comparison of different variants. Sup-only: our method uses only labeled data. Feature: Feature-level perturbation consistency. Image: Image-level strong-to-weak consistency. Ours: Gate-guided two-level perturbation consistency.
Existing SSCD methods can be broadly divided into three categories: adversarial learning-based, pseudo-label-based, and consistency regularization-based. AdvEnt [5], SemiCDNet [6], and SALCD [7] are typical adversarial learning-based methods that employ alternative optimization strategies to improve the representation learning of their respective models. Pseudo-label-based methods, such as RC-CD [8], SemiSiROC [9], and DCF-CFe [10], emphasize the improvement of pseudo-label quality and apply some techniques to enhance feature distinctiveness. Consistency regularization-based methods, including SemiSANet [11], SemiBuildingChange [12], and SemiPTCD [13], enforce that perturbed images or features produce identical outputs as the original inputs [14].
Recently, the consistency regularization framework has gained increasing attention in SSCD due to its simplicity and stability. However, existing methods mainly focus on modifying network architectures or incorporating supplementary information based on weak-to-strong consistency. For instance, SemiSANet builds a siamese nested UNet with graph attention, MTCNet [15] augments the change labels with additional segmentation labels, and ST-RCL [16] combines self-training and consistency learning. Meanwhile, feature-level consistency remains relatively simplistic and underdeveloped, such as UniMatch [17] applying dropout for feature consistency. Thus, there exists substantial potential in integrating image-level and feature-level consistency regularization to enhance SSCD’s performance and robustness.
@@
As our research indicates, simply employing a basic SSCD architecture and applying perturbations only at the image or feature level fails to achieve satisfactory results. Fig. 1 shows the results of networks employing perturbations at different perspectives, using $5\%$ labeled data and $95\%$ unlabeled data on two CD datasets [18, 6]. The visual cases and performance comparison strongly prove this viewpoint. Although perturbations contribute to enhancing model robustness, relying solely on image perturbations may affect pixel integrity and segmentation, and exclusive dependence on feature perturbations might struggle with image diversity.
@@
In this paper, we explore consistency regularization at a broader scope and propose a Gate-guided Two-level Perturbation Consistency regularization-based SSCD method (GTPC-SSCD). By learning both image-level strong-to-weak consistency and feature-level perturbation consistency, GTPC-SSCD can exploit the unlabeled data more effectively, thereby achieving a more thorough comprehension of the data characteristics. Nevertheless, performing the same treatment on all samples constrains the potential efficacy of perturbation, particularly evident in remote sensing change detection scenarios. To alleviate this issue, we develop a hardness analysis-based gating mechanism. It is designed to distinguish the difficulty of training samples and determine the application of feature perturbation accordingly. We carry out comparative experiments on six CD datasets with seven state-of-the-art methods, and the proposed GTPC-SSCD achieves the best performance.
Our contributions are summarized as follows:
@@
  • We separately analyze image-level and feature-level perturbation implementations in SSCD and combine them with adaptive strategies to improve performance.
@@
  • We propose GTPC-SSCD, a Gate-guided Two-level Perturbation Consistency regularization-based SSCD method. And we design a hardness analysis-based gating mechanism to assess the difficulty of samples and determine the necessity of applying feature perturbations.
@@
  • We conduct extensive experiments on six benchmark CD datasets and our method exhibits higher accuracy than other seven state-of-the-art methods.
@@ II Methodology
Semi-supervised change detection (SSCD) utilizes a limited amount of labeled data and a large amount of unlabeled data to train a change detection network for accurate change maps generation. The training set consists of two subsets, a labeled set and an unlabeled set. The labeled set can be represented as $\mathcal{D}_{l}=\{{(\mathbf{X}^{l}_{A,i},\mathbf{X}^{l}_{\mathrm{B},i}),\mathbf{Y}^{l}_{i}}\}^{M}_{i=1}$, where $(\mathbf{X}^{l}_{A,i},\mathbf{X}^{l}_{B,i})$ denotes the $i$-th labeled image pair, $\mathbf{X}^{l}_{A,i}$ is the pre-change image, $\mathbf{X}^{l}_{B,i}$ is the post-change image, and $\mathbf{Y}^{l}_{i}$ is the corresponding ground truth. Let $\mathcal{D}_{u}=\{{(\mathbf{X}^{u}_{A,i},\mathbf{X}^{u}_{B,i})}\}^{N}_{i=1}$ denotes the unlabeled set. $(\mathbf{X}^{u}_{A,i},\mathbf{X}^{u}_{B,i})$ is the $i$-th unlabeled image pair. $M$ and $N$ indicate the number of labeled and unlabeled image pairs, respectively. In most cases, we have $N>>M$. In the following sections, we will detail the proposed GTPC-SSCD.
II-A Our GTPC-SSCD method
Our method consists of a supervised training part and an unsupervised training part.
@@ II-A Our GTPC-SSCD method
II-A1 Supervised training part
We use the labeled set $\mathcal{D}_{l}$ to train the CD network $\phi$. The network processes weakly augmented image pairs to generate change map $\mathbf{P}^{l}$. We utilize the standard cross-entropy (CE) loss as supervision. The supervised loss $\mathcal{L}_{s}$ is formulated as:
$\mathcal{L}_{s}=\mathcal{L}_{CE}(\mathbf{P}^{l},\mathbf{Y}^{l}),$ (1)
where $\mathcal{L}_{CE}$ represents the CE loss, which measures the dissimilarity between the predicted change map $\mathbf{P}^{l}$ and the ground truth $\mathbf{Y}^{l}$.
II-A2 Unsupervised training part
As illustrated in Fig. 2, we apply two-level perturbations to train $\phi$ using the unlabeled set $\mathcal{D}_{u}$. Specifically, we enforce strong-to-weak consistency at the image level and maintain perturbation consistency at the feature level. The former enhances the model’s robustness against various data transformations like illumination and season changes, while the latter sharpens the model’s generalization and interference resistance.
Image-level strong-to-weak consistency learning. Drawing inspiration from ConMatch [19], we expand the quantity of strong augmentation branches, and enforce the strong-to-weak consistency by aligning strong and weak outputs. It promotes a more efficient exploitation of the latent potential within unlabeled data, strengthening the model’s adaptability across diverse data scenarios and enhancing its robustness. Specifically, we apply two independent strong augmentations on $(\mathbf{X}^{uw}_{A},\mathbf{X}^{uw}_{B})$, resulting in $(\mathbf{X}^{us_{1}}_{A},\mathbf{X}^{us_{1}}_{B})$ and $(\mathbf{X}^{us_{2}}_{A},\mathbf{X}^{us_{2}}_{B})$. They are fed into $\phi$ and generate three change maps $\mathbf{P}^{uw}$, $\mathbf{P}^{us_{1}}$ and $\mathbf{P}^{us_{2}}$.
@@ II-A2 Unsupervised training part
$\hat{\mathbf{Y}}^{uw}=\begin{cases}1,&if\ \mathbf{P}^{uw}>\tau\\ 0,&else\end{cases}$ (2)
where $\tau$ = 0.95 is a confidence threshold. The image-level consistency loss $\mathcal{L}_{ui}$ is expressed as:
$\mathcal{L}_{ui}=\frac{1}{2}(\mathcal{L}_{CE}(\mathbf{P}^{us_{1}},\hat{\mathbf{Y}}^{uw})+\mathcal{L}_{CE}(\mathbf{P}^{us_{2}},\hat{\mathbf{Y}}^{uw})).$ (3)
Here, weak augmentations consist of random resizing, cropping, and horizontal flipping. Strong augmentations include random color jittering, Gaussian blur, and CutMix.
Feature-level perturbation consistency learning. To establish feature-level consistency, we implement several different feature perturbations on change features and ensure the perturbed outputs align with the originals. The process of generating change maps $\mathbf{P}^{uf}_{1:K}$ is formulated as:
$\mathbf{P}^{uf}_{1:K}=f^{{}^{\prime}}_{d,1:K}(\mathbf{D}^{u}_{4},\mathbf{D}^{u^{\prime}}_{1,1:K}),$ (4)
where $\mathbf{D}^{u^{\prime}}_{1,1:K}=\{\mathbf{D}^{u^{\prime}}_{1,k}\}^{K}_{k=1}$, represents the difference features after perturbation. $f^{{}^{\prime}}_{d,1:K}=\{f^{{}^{\prime}}_{d,k}\}^{K}_{k=1}$, denote the auxiliary decoders. $\mathbf{P}^{uf}_{1:K}=\{\mathbf{P}^{uf}_{k}\}^{K}_{k=1}$. The feature-level consistency loss $\mathcal{L}_{uf}$ is expressed as:
$\mathcal{L}_{uf}=\frac{1}{K}\sum^{K}_{k=1}\mathcal{L}_{CE}(\mathbf{P}^{uf}_{k},\hat{\mathbf{Y}}^{uw}),$ (5)
where $K$ is the total number of the auxiliary decoders. We adopt seven types of feature perturbations [20], i.e., feature noise, feature dropout, object masking, context masking, guided cutout, Intermediate VAT, and random dropout.
Hardness analysis-based gating mechanism. To boost the model’s adaptability to samples with varying complexity, we develop a hardness-based gating mechanism. It utilizes an additional decoder to generate a change map $\mathbf{P}^{uw^{\prime}}_{i}$. Then, an Intersection over Union (IoU) comparison is performed between $\mathbf{P}^{uw^{\prime}}_{i}$ and $\mathbf{P}^{uw}_{i}$ to quantitatively assess the complexity of each sample. Utilizing the gating mechanism, we can select challenging samples and apply various perturbations on $\mathbf{D}^{u}_{1}$ of these selected samples to obtain perturbed features. The whole process can be formulated as:
$iou_{i}=\mathrm{IoU}(\mathbf{P}^{uw^{\prime}}_{i},\mathbf{P}^{uw}_{i}),$ (6)
@@
where $iou_{i}$ denotes the IoU score of the $i$-th sample. $\mathrm{Mid}(iou)$ is calculated as the median value, serving as a stable and reliable ratio. $\mathrm{FP}_{1:K}(\cdot)$ denotes $K$ different perturbations to the input features. $\mathbf{D}^{u^{\prime}}_{1,1:K}=\{\mathbf{D}^{u^{\prime}}_{1,k}\}^{K}_{k=1}$, is the set of $\mathbf{D}^{u^{\prime}}_{1,k}$. Through this differential sample processing strategy, the model can effectively harness the potential of unlabeled data, enhancing its overall performance.
II-A3 Loss function
The total loss consists of the supervised loss $\mathcal{L}_{s}$, the image-level consistency loss $\mathcal{L}_{ui}$ and the feature-level consistency loss $\mathcal{L}_{uf}$. It is formulated as:
@@ II-B1 Change detection network
Our CD network $\phi$ consists of two components: a difference feature generator $f_{g}$ and a decoder $f_{d}$. The generation of change map $\mathbf{P}$ is expressed by:
@@ II-B2 Super-parameters
The experiments are implemented using PyTorch on an NVIDIA RTX2080Ti GPU. Our model adopts the SGD optimizer with a learning rate of 0.02, a momentum of 0.9, and a weight decay of 1e-4. All models are trained for 80 epochs with a batch size of 4 for both labeled and unlabeled data.
III Experiment
III-A Experimental Setup
Baselines. To validate the effectiveness and superiority of our proposed method, we compare it with seven existing state-of-the-art SSCD methods, including AdvEnt [5], s4GAN [23], SemiCDNet [6], SemiCD [24], RC-CD [8], SemiPTCD [13], and UniMatch [17]. Notably, the first three are adversarial learning-based methods, RC-CD is a pseudo-label-based method, and the remaining three are based on consistency regularization. Sup-only refers to our method trained only on labeled data. All methods are conducted with PyTorch and trained on the same training sets.